OpenStudy (anonymous):

Solve log 2x^2 + 2 = 6. Round to the nearest thousandth if necessary.

OpenStudy (anonymous):

I though i'd reopen a new one. Was afk for a long time

Nnesha (nnesha):

so parentheses isn't there ? right ?

Nnesha (nnesha):

you can tag me 'n the last post :=)

OpenStudy (anonymous):

nope no parenthesis look just like this \[2x^2 + 2 =6\]

Nnesha (nnesha):

alright so you work is correct!

Nnesha (nnesha):

ur*

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @lostinprague I subtract 2 from both sides log10 (2x^2) + 2 -2 = 6-2 log10 (2x^2)=4 I then used a = logb (b^a) so... 4= log10 (10^4) log10 (2x^2) = log10 (10^4) then all this other stuff/ not sure I want to type it all out lol \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

okay so it was 50squareroot 2 or 70.711 then? I think thats what I had down

OpenStudy (anonymous):

yeah thats my work

OpenStudy (anonymous):

thanks for helping me :)

Nnesha (nnesha):

yes right np :=)